The rejection region for a 5% level test of H0 : μ ≥ 10 versus H1 : μ < 10 is X¯ < 7.8. Find the rejection region for a 1% level test.
Answer #1

solution:

the given information as follows:

null and alternative hypothesis:

H_0:\mu \geq 10

H_1:\mu < 10

the test is a one tailed test

g= 0,05

rejecttion region = \bar X<7.8 (which mean if mean is less than 7.8 null will be rejected at 5%)

so critical value of z at g= 0,05 for left tail from z table = -1.645

so finding the standard error as follows:

z = \frac{\bar X-\mu}{se}

Se=\frac{\bar X-\mu}{z}=\frac{7.8-10}{-1.645}=1.3374

now we have to find the mean score for which null will be rejected at 1%

so α = 0,01

critical value of z = z_c=z_{0.01}=-2.33

z = \frac{\bar X-\mu}{se}

\bar X=(z * \sigma)+\mu=(-2.33*1.3374)+10=6.88

so , rejection region at 1% = \bar X < 6.88